Operation Summary for September 11, 2009 - September 14, 2009 |
Friday September 11 Startup continued. Main Injector (MI) personnel conducted beam studies. The Antiproton Source (Pbar) and the Recycler used protons to tune their machines. During the day and evening shifts, Operations received permission to run beam in the TeV early on the day shift and iterations of turn-on and quench recovery followed. MI-8, MI, Recycler, NuMI, Pbar, and the Tevatron (TeV) all conducted beam studies. They encountered and resolved many machine problems. Beam was delivered to MiniBooNE when available. Pbar began stacking and continued machine tune up work. NuMI took beam for aperture, target, and horn scans. Saturday September 12 The midnight shift began with Operations monitoring the stash (13.7E10) and the stack (18.0E10), and with startup operations and studies in progress for the TeV, MI, Recycler, Switchyard, NuMI, and MiniBooNE. During the midnight shift, a Pbar Debuncher quadrupole power supply (DQ719) tripped off requiring attention from Pbar personnel. A quench at TeV sector F1 ended beam studies for the night and a subsequent quench at sector F3 delayed an extended dry squeeze. TeV studiers arrived at the end of the midnight shift to continue beam studies. MI and Recycler personnel also conducted studies. During the day shift, NuMI studiers finished up and operators prepared the enclosures for access. Startup studies and commissioning activities continued. During the evening shift, NuMI studies and target scans finished up early in the evening and operators prepared MI-65 prepared for access. Operations successfully transferred antiprotons to the Recycler; cooling studies followed. Pbar stacked during the shift. A MI RF station (MIRF13) tripped off several times due to a shunt regulator over-current fault. A Debuncher power supply (D:H914) faulted off with a phase unbalanced trip; Operations called in personnel to repair it. The TeV quenched while parsing the squeeze. Operations delivered beam to the Pbar and MiniBooNE when available. Sunday September 13 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Recycler, Pbar, and TeV beam studies and machine commissioning continued. Pbar personnel were called in to resolve power supply issues. A MI Lambertson (LAM52) tripped off on magnet Klixon fault and would not reset. Operators investigated and repaired a bad connection. Operations completed one transfer of antiprotons to the Recycler for beam studies. TeV studiers parsed their way through the squeeze until a quench at D0 occurred due in part to a TeV Sector D0 quadrupole power supply (D0Q6). Operations called in EE Support personnel to investigate; they swapped power supply regulating hardware with another power supply (B0Q6). TeV checkout and turn on followed. During the day shift, Recycler cooling node problems required expert assistance, which delayed beam studies. NuMI accessed the absorber area following beam studies. Beam studies resumed after the access until the horn power supply tripped its wall breaker. EE Support personnel came in and investigated the problem. During the evening shift, a Booster RF station (BRF11) tripped off with a power amplifier LCW flow trip indication. The station was left off. The Pbar lithium lens tripped off several times with a false LCW flow trip indication. MI collimator studies took place at various times throughout the shift. Beam was delivered to Pbar and MiniBooNE when available. Monday September 14 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, EE Support personnel reported that no significant issues were found with the NuMI horn and it was returned to service. The NuMI extraction kicker high voltage power supply failed and delayed the continuation of beam studies for a few hours. Operations prepared MI-65 for access. Beam was delivered to the Pbar despite continuing problems with the low conductivity interlock for the Pbar lens. Pbar target personnel arrived to investigate. Recycler beam studies were conducted for most of the shift. Beam was delivered to MiniBooNE when available. Operations reported the loss of the bypass loop for the TeV ramp on two occasions prompting operators to dump the ramp. In between loop trips the TeV quenched at sector B3 when a heater firing unit discharged prematurely; the unit was successfully changed out. The TeV ramp was left off after the second bypass loop trip so EE Support personnel could swap out the ground fault detector. The Plans for Monday Continue with startup. MI might access on Tuesday to realign magnets. |
|
More Information Comments and Suggestions |
| last modified 09/14/2009 email Fermilab |
| Security, Privacy, Legal | |